A 4.00-g copper coin at 17.5°C drops 55.0 m to the ground.

(a) Assuming 50.0% of the change in gravitational potential energy of the coin–Earth system goes into increasing the internal energy of the coin, determine the coin's final temperature. °C

(b) Does the result depend on the mass of the coin?

Yes or No  


Explain your answer.
Answer #1

given
m = 4 g = 0.004 kg
Ti = 17.5 C
h = 55 m

we know, specific heat of copper, C = 385 J/(kg K)

a) heat gained by coin = 50% of loss of potential energy

m*C*(Tf - Ti) = 0.5*m*g*h

Tf - Ti = 0.5*g*h/C

Tf = Ti + 0.5*g*h/C

= 17.5 + 0.5*9.8*55/385

= 18.2 degrees celsius

b) No

because, final temperature, Tf = Ti + 0.5*g*h/C

in final temperature we don't have mass term.